Jane is deciding whether to go to school for 8 weeks this summer. The cost of tuition and textbooks is $1,700 and housing and other expenses will cost her $600. If she does not go to school, she will live in her parents house for free and they will cover her food and other expenses for her. Also, if Jane does not go to summer school she could work fulltime. But the best job she can get pays only $600 per week, and Jane would only agree to give up her free time for no less than $750 per week. However, if she goes to summer school, she1l have to spend 40 hours a week attending classes and studying a) b) c) What will the summer school cost Jane in terms of money spent? (5 marks) What are the opportunity costs of going to summer school that Jane does not pay explicitly? Explain What is Janes total opportunity cost of going to school this summer? Explain your answer. (5 marks) d) Suppose that if Jane does not go to summer school, she will eventually take the dasses anyway What is Janes marginal benefit of going to summer school? (5 marks) e) Suppose Jane decides to go to school in the summer. Explain her decision using the concepts of marginal cost and marginal benefit. (5 marks)|
Answer #1

A) Cost of summer school(accounting/explicit costs)= Housing and other expenses+ Tuition fee= $1700+$600=$2300

B) The job she could have done i.e. the money that could have been made working instead of going to school and the value she puts on her free time which is $750

C) Jane's total opportunity cost of going to school this summer will be $600*8=$4800. The next best alternative she is getting instead of attending a summer school for 8 weeks is $600. So she is foregoing $600 for 8 weeks totalling upto $4800.

D) Jane's marginal benefit of going to summer school should be at least $150 per week, that is the amount of difference in the opportunity cost($600) and what she values her time($750)

E) Total cost(implicit+explicit)= $2300+$4800=$7100

Marginal cost per week=$7100/7=$1014.28

Total benefit(explicit+implicit)= Saving on other expenses($600)+value of free time on studying($750*8)=$6600

Marginal benefit per week=$6600/942.85

MC>MB. Hence, she will not attend summer school.